OpenStudy (anonymous):

Can someone please help me with this, like explain how to do it cause I have no idea :/ Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2). A. Y= -1/3x -6 B. Y= -1/3x-14 C. Y=3x-6 D. Y=3x-14

OpenStudy (anonymous):

parallel lines have the same slope. so the first thing you would have to do is find the equation that passes through those points by using: point slope form which is y-y1=m(x-x1) m=slope and you use (4,-2) y+2=3(x-4) y+2=3x-12 y=3x-14 soi the answer is D.

OpenStudy (anonymous):

Thank you so much I still don't get it :/ but it's okay I'll try to figure it out